# Ramsey Numbers in Kneser Graphs

### v1

@@ 3.2 Lower bounds
The lower bound in Proposition 10 takes advantage of the chromatic number of Kneser graphs. Lovász  [5] proved that $\chi(\textrm{KG}(n,r))=n-2r+2$, which was one of the earliest applications of topological methods to combinatorics.
@@
Let $n=R^{\textrm{KG}}_{r}(s,t)$ and set $G=\textrm{KG}(n,r)$. Suppose $\chi(G)=n-2r+2<R(s,t)$ and let $c:V(G)\to[n-2r+2]$ be a proper vertex-coloring of $G$. Consider the complete graph $G^{\prime}$ on n2r+2n-2r+2 vertices with $V(G^{\prime})=[n-2r+2]$ and let $C^{\prime}:E(G^{\prime})\to\{R,B\}$ be a red/blue edge-coloring of $G^{\prime}$ avoiding a red copy of $K_{s}$ and a blue copy of $K_{t}$.
Now, consider the following red/blue edge-coloring of $G$: For $uv\in E(G)$, color $uv$ with $C^{\prime}(c(u),c(v))$. We claim that this coloring has no red copy of $K_{s}$ or blue copy of $K_{t}$. Suppose there is a red copy of $K_{s}$ with vertices $\{v_{1},v_{2},\dots,v_{s}\}$. Observe that for all $i\neq j$, $c(v_{i})\neq c(v_{j})$ since $c$ is a proper vertex-coloring and $v_{i}v_{j}\in E(G)$. But now, the color of $v_{i}v_{j}$ is $C^{\prime}(c(v_{i}),c(v_{j}))=R$. This implies that the graph induced by the vertices $\{c(v_{1}),c(v_{2}),\dots,c(v_{s})\}$ is a red copy of $K_{s}$, a contradiction to our choice of $C^{\prime}$. Therefore, $n-2r+2\geq R(s,t)$. ∎
@@
Observe that the induced subgraph $G[Z]$ of $G$ on $Z$ is triangle-free. Also, $X$ and $Y$ are independent sets with no triangles contained in $G[X\cup Y]$. Therefore, the only triangles will have either two vertices from $Z$ and one from $X\cup Y$ or one vertex from each of $X,Y,Z$. Setting $c(E(G[Z]))=B$ and $c(E(X,Y))=B$ with all other edges colored $R$ (see Figure 1), we obtain a red/blue coloring with no monochromatic triangles. ∎
@@ 3.3 Computationally-obtained bounds
The proof of Theorem 12 follows because our program found a good coloring of $\textrm{KG}(8,2)$ (this coloring is given explicitly in Appendix A), but declared that the graphs $\textrm{KG}(9,2)$ and $\textrm{KG}(13,3)$ do not have good colorings.
Our code is given in Appendix B.
@@ 4 Pálvölgyi’s problem
We begin with a general discussion on monochromatic induced Kneser subgraphs, which will be useful in making progress on Problem 13.
Let $r>1$ be an integer and let $\textrm{KG}(s,r)$ and $\textrm{KG}(t,r)$ be Kneser graphs. Define the induced $r$-Kneser Ramsey number $R^{\textrm{KG}}_{r,ind}(s,t)$ to be the minimum $n$ (if it exists) such that every red/blue edge-coloring of $\textrm{KG}(n,r)$ contains an induced red $\textrm{KG}(s,r)$ or an induced blue $\textrm{KG}(t,r)$.
Theorem 18.
For $r\geq 2$, $R^{\textrm{KG}}_{r,ind}(s,t)$ does not exists; that is, there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no red induced $\textrm{KG}(s,r)$ or blue induced $\textrm{KG}(t,r)$ .
In the proof of Theorem 18 we use the set-system theorem of Bollobás.
@@
Proof of Theorem 18.
We first show $R^{\textrm{KG}}_{r,ind}(2r,2r)$ does not exist. Observe that $\textrm{KG}(2r,r)$ is a matching of size $\frac{1}{2}\binom{2r}{r}$.
Now, in any $\textrm{KG}(n,r)$ with $n\geq 2r$, we claim that every maximum induced matching $M$ has vertex set $\binom{S}{r}$, where $|S|=2r$. To this end, let $M$ be an induced matching with edge set $E=\{(X_{i},Y_{i})\colon 1\leq i\leq m\}$. Note that we must have $X_{i}\cap Y_{i}=\emptyset$ for all $i$, and $X_{i}\cap Y_{j}\neq\emptyset$ for all $i\neq j$ in order for $M$ to be an induced matching. Let $A_{i}=X_{i}$ for $1\leq i\leq m$, $A_{i}=Y_{i-m}$ for $m+1\leq i\leq 2m$, and let $B_{i}=Y_{i}$ for $1\leq i\leq m$, $B_{i}=X_{i-m}$ for $m+1\leq i\leq 2m$. Then the set pair system $(A_{i},B_{i})$, $1\leq i\leq 2m$ satisfies $A_{i}\cap B_{i}=\emptyset$ for all $i$ and $A_{i}\cap B_{j}\neq\emptyset$ for $i\neq j$.
Thus, we may invoke Theorem 19, with $a=b=r$, and
$2\cdot|E(M)|=2m\leq\binom{2r}{r}$.
In fact, since we know there exists an induced matching of size $\frac{1}{2}\binom{2r}{r}$, equality can be achieved, and there is a set $S$ such that $|S|=2r$ where $A_{i}\subseteq S$ and $B_{i}=S\setminus A_{i}$ for each $i$. Hence, one edge in a maximum induced matching $M$ uniquely determines the $2r$ elements of the ground set which can be used to form $M$.
Since a maximum induced matching $M$ can be uniquely determined by a single edge in $\textrm{KG}(n,r)$, this implies any two maximum induced matchings, i.e. matchings isomorphic to $\textrm{KG}(2r,r)$, are disjoint, and we can color each maximum induced matching bichromatically. Thus, we have a coloring of any $\textrm{KG}(n,r)$ with no monochromatic $\textrm{KG}(2r,r)$; i.e. $R^{\textrm{KG}}_{r,ind}(2r,2r)$ does not exist.
@@
Finally, observe that for any $s,t\geq 2r$, $\textrm{KG}(s,r)$ and $\textrm{KG}(t,r)$ each contain a subgraph isomorphic to $\textrm{KG}(2r,r)$. Hence, since there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no induced monochromatic $\textrm{KG}(2r,r)$ for any $n\geq 2r$, in the same edge-coloring there are no monochromatic induced $\textrm{KG}(s,r)$ nor monochromatic induced $\textrm{KG}(t,r)$. ∎
Proof of Theorem 14.
Set $n=R_{r}(3r,3r)-1$. We aim to show that there is a red/blue coloring $c$ of $G=\textrm{KG}(n,r)$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle. We begin by partitioning $V(G)$ into two sets. Consider a red/blue edge coloring $c^{\prime}$ of $G^{\prime}=K^{r}_{n}$ containing no monochromatic $K^{r}_{3r}$. Observe that $E(G^{\prime})=V(G)$ and there is a clear bijection between the two sets. Let us partition $V(G)$ into two sets $P_{1},P_{2}$ as follows: For $e=\{v_{1},v_{2},\dots,v_{r}\}\in E(G^{\prime})$, if $c(e)=R$, add $\{v_{1},v_{2},\dots,v_{r}\}$ to $P_{1}$. Otherwise, add $\{v_{1},v_{2},\dots,v_{r}\}$ to $P_{2}$.
@@
Observe that by the choice of $c^{\prime}$, neither $G[P_{1}]$ nor $G[P_{2}]$ contain an induced copy of $\textrm{KG}(3r,r)$ (since otherwise, $c^{\prime}$ would contain a monochromatic $K^{r}_{3r}$). Therefore, for any edge $e$ within $G[P_{1}]$ or $G[P_{2}]$, set $c(e)=R$ and otherwise, set $c(e)=B$. Since the blue subgraph is a bipartite graph, it is triangle-free. And so, $c$ is a red/blue edge coloring of $G$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle. ∎
@@ 5 Concluding remark
Appendix A: A good coloring of $\textrm{KG}(8,2)$.
Here, we provide a coloring of $\textrm{KG}(8,2)$ with no monochromatic triangle.
@@
((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,4),(2,6)),((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,4),(5,6)),((1,4),(6,8)),((1,5),(2,3)),((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,5),(6,7)),((1,5),(6,8)),((1,5),(7,8)),((1,6),(2,3)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(2,3)),((1,7),(2,5)),((1,7),(2,8)),((1,7),(3,5)),((1,7),(3,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(2,3)),((1,8),(2,4)),((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),((2,4),(3,6)),((2,4),(3,8)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(3,6)),((2,5),(3,7)),((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(3,4)),((2,6),(3,5)),((2,6),(4,5)),((2,6),(5,7)),((2,6),(5,8)),((2,7),(3,5)),((2,7),(3,8)),((2,7),(4,5)),((2,7),(4,8)),((2,7),(5,8)),((2,8),(3,4)),((2,8),(3,7)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),((3,8),(6,7)),((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),\\ ((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,4),(2,6)),\\ ((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,4),(5,6)),((1,4),(6,8)),((1,5),(2,3)),\\ ((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,5),(6,7)),((1,5),(6,8)),\\ ((1,5),(7,8)),((1,6),(2,3)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),\\ ((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(2,3)),((1,7),(2,5)),((1,7),(2,8)),\\ ((1,7),(3,5)),((1,7),(3,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(2,3)),((1,8),(2,4)),\\ ((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),\\ ((2,4),(3,6)),((2,4),(3,8)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(3,6)),((2,5),(3,7)),\\ ((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(3,4)),((2,6),(3,5)),((2,6),(4,5)),\\ ((2,6),(5,7)),((2,6),(5,8)),((2,7),(3,5)),((2,7),(3,8)),((2,7),(4,5)),((2,7),(4,8)),\\ ((2,7),(5,8)),((2,8),(3,4)),((2,8),(3,7)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),\\ ((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),\\ ((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),\\ ((3,8),(6,7)),((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),\\ ((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),\\ ((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))

### v2

@@ 3.2 Lower bounds
For the lower bound in Proposition 10 we employ the chromatic number of Kneser graphs. It is easy to verify that $\chi(\textrm{KG}(n,r))\geq n-2r+2$ by exhibiting a explicit proper coloring, and this is the bound we need for the proof. Lovász [5] proved that $\chi(\textrm{KG}(n,r))=n-2r+2$, which was one of the earliest applications of topological methods to combinatorics.
@@
For $r\geq 1$ and $s,t\geq 2$, we wish to show $R^{\textrm{KG}}_{r}(s,t)\geq R(s,t)+2r-2.$ Let $n=R^{\textrm{KG}}_{r}(s,t)$ and set $G=\textrm{KG}(n,r)$. Assume for contradiction that $R(s,t)>n-2r+2=\chi(G)$ and let $c:V(G)\to[n-2r+2]$ be a proper vertex-coloring of $G$. Consider the complete graph $G^{\prime}$ on vertex set [n2r+2][n-2r+2] and let $c^{\prime}:E(G^{\prime})\to\{R,B\}$ be a red/blue edge-coloring of $G^{\prime}$ avoiding a red copy of $K_{s}$ and a blue copy of $K_{t}$. Note that $c^{\prime}$ exists because of our negation assumption.
Now, define a red/blue edge-coloring of $f:E(G)\to\{R,B\}$ by
$f(uv)=c^{\prime}(c(u),c(v))$
for $uv\in E(G)$. We claim that $f$ has no red copy of $K_{s}$ or blue copy of $K_{t}$. Suppose there is a red copy of $K_{s}$ with vertices $\{v_{1},v_{2},\dots,v_{s}\}$. Then for $i\neq j$ we have
$f(v_{i}v_{j})=c^{\prime}(c(v_{i}),c(v_{j}))=R.$
Observe that when $i\neq j$, $c(v_{i})\neq c(v_{j})$ because $c$ is a proper vertex-coloring of $G$ and $v_{i}v_{j}\in E(G)$. This implies that the subgraph of $G^{\prime}$ induced on the vertices $\{c(v_{1}),c(v_{2}),\dots,c(v_{s})\}$ is a red copy of $K_{s}$, a contradiction to our choice of $c^{\prime}$. A similar argument shows that $f$ has no blue copy of $K_{t}$. ∎
@@
Observe that the induced subgraph $G[Z]$ of $G$ on vertex set $Z$ is triangle-free. Also, $X$ and $Y$ are independent sets with no triangles contained in $G[X\cup Y]$. Therefore, the only triangles will have either two vertices from $Z$ and one from $X\cup Y$ or one vertex from each of $X,Y,Z$. Setting $c(E(G[Z]))=B$ and $c(E(X,Y))=B$ with all other edges colored $R$ (see Figure 1), we obtain a red/blue coloring with no monochromatic triangles. ∎
@@ 3.3 Computationally-obtained bounds
The proof of Theorem 12 for $R^{\textrm{KG}}_{2}(3,3)$ follows because our program found a good coloring of $\textrm{KG}(8,2)$, but declared that the graph $\textrm{KG}(9,2)$ does not have a good coloring. The bounds for the other Kneser Ramsey numbers in Theorem 12 were obtained in similar ways.
The explicit colorings of $\textrm{KG}(8,2)$ with no monochromatic $K_{3}$, and of $\textrm{KG}(11,2)$ with no red $K_{4}$ or blue $K_{3}$ are given in Appendix A. Our code is given in Appendix B.
@@ 4 Pálvölgyi’s problem
4.1 Induced Kneser graphs
We start with a discussion on induced graphs of the Kneser graph. Recall that given integer $n\geq s\geq r$, we defined an induced $\textrm{KG}(s,r)$ as a subgraph $H$ of the graph $\textrm{KG}(n,r)$ induced on a vertex set $S\choose r$}, where $S\subset[n]$ is a ground set of size $s$.
We first show that every induced graph $H$ of $\textrm{KG}(n,r)$ isomorphic to $\textrm{KG}(2r,r)$ is, in fact, an induced $\textrm{KG}(2r,r)$:
Proposition 17.
Suppose $H$ is an induced subgraph of $G=\textrm{KG}(n,r)$ isomorphic to $\textrm{KG}(2r,r)$. Then $H$ is an induced $\textrm{KG}(2r,r)$, that is, there exists a ground set $S\subset[n]$ of size $2r$ such that $H=G[{S\choose r}$].
The proof will follow from the set-system theorem of Bollobás.
@@ 4.1 Induced Kneser graphs
Proof of Proposition 17.
Observe that $\textrm{KG}(2r,r)$ is a matching of size $\frac{1}{2}\binom{2r}{r}$. We claim that every induced matching in $\textrm{KG}(n,r)$ of size $\frac{1}{2}\binom{2r}{r}$ has vertex set $\binom{S}{r}$, for some $S\subset[n]$ and $|S|=2r$. Indeed, let $M=\{(X_{i},Y_{i})\colon 1\leq i\leq m\}$ be an induced matching of $\textrm{KG}(n,r)$. Then $X_{i}\cap Y_{i}=\emptyset$ for all $i\in[m]$, and $X_{i}\cap Y_{j}\neq\emptyset$ for all $i\neq j$.
Set
[…]
and
[…]
Then the set pair system $\{(A_{i},B_{i})\mid 1\leq i\leq 2m\}$ satisfies the conditions of Theorem 18. Therefore, by the theorem,
$2|M|=2m\leq\binom{2r}{r}$,
and moreover, equality is achieved when the sets $A_{i}$ are all the $r$ subsets of the set $S=A_{1}\cup B_{1}$, where $B_{i}=S\setminus A_{i}$ for all $i$. This implies that $H$ is a induced $KG(2r,r)$ on ground set $S$, as claimed. ∎
We return now to Pálvölgyi’s problem. In Problem 13 one would like to find a red/blue edge-coloring of $G=\textrm{KG}(n,r)$ with no red induced $\textrm{KG}(3r,r)$ and no blue triangle. We first show the existence of a red/blue edge-coloring of $G$ with no monochromatic induced Kneser graphs $\textrm{KG}(s,r)$ for all $n\geq s\geq 2r$.
Proposition 19.
@@
Let $n\geq 2r\geq 4$ be integers. Then there exists a red/blue edge-coloring of $\textrm{KG}(n,r)$ with no monochromatic induced $\textrm{KG}(s,r)$ for all $n\geq s\geq 2r$.
Proof.
We say that an edge $XY\in\textrm{KG}(n,r)$ is interlacing, where $X,Y\in{[n]\choose r}$ are disjoint sets, if there exist $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$ such that either $y<x<y^{\prime}$ or $x<y<x^{\prime}$. Otherwise, we say that $XY$ is non-interlacing. Now color the interlacing edges of $\textrm{KG}(n,r)$ red, and the non-interlacing edges blue.
We claim that in this coloring there is no monochromatic $\textrm{KG}(s,r)$ for any $n\geq s\geq 2r$. Indeed, suppose $H$ is an induced $\textrm{KG}(s,r)$ on ground set $S=\{z_{1}<z_{2}<\dots<z_{s}\}\subset[n]$ where $s\geq 2r\geq 4$. Then the edge $XY\in H$, where $X=\{z_{1},\dots z_{r}\}$ and $Y=\{z_{r+1},\dots z_{2r}\}$ is non-interlacing, while the edge $X^{\prime}Y^{\prime}\in H$, where $X^{\prime}=\{z_{2},\dots z_{r+1}\}$ and $Y^{\prime}=\{z_{1},z_{r+2}\dots z_{2r}\}$ is interlacing, showing that $H$ is not monochromatic. ∎
The proof shows that an induced $\textrm{KG}(s,r)$ is not monochromatic because it contains an induced $\textrm{KG}(2r,r)$, which is already not monochromatic under the edge coloring in the proof. Therefore, by Proposition 17 we can state Proposition 19 in a slightly stronger manner:
Corollary 20.
Let $n\geq 2r\geq 4$ be integers. There exists a red/blue edge-coloring of $G=\textrm{KG}(n,r)$ such that for all $n\geq s\geq 2r$, if $H$ is an induced subgraph of $G$ isomorphic to $\textrm{KG}(s,r)$, then $H$ is not monochromatic.
Note that Problem 13 is not resolved by Corollary 20 because a triangle is an induced Kneser graph only if $r=1$. In the next subsection we prove Theorem 14, which gives a lower bound for Problem 13.
4.2 Proof of Theorem 14
Set $n=R_{r}(3r,3r)-1$, where $R_{r}(3r,3r)$ is the minimum integer $m$ such that in any red/blue coloring of the edges of the complete $r$-uniform hypergraph on $m$ vertices $K^{r}_{m}$ there is a monochromatic $K^{r}_{3r}$. We would like to show that there is a red/blue coloring $c$ of $G=\textrm{KG}(n,r)$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle.
Let $c^{\prime}$ be a red/blue edge coloring of the graph $G^{\prime}=K^{r}_{n}$ containing no monochromatic $K^{r}_{3r}$ ($c^{\prime}$ exists by the choice of $n$). We partition $E(G^{\prime})=V(G)$ into two sets $P_{1},P_{2}$ as follows: for $e\in E(G^{\prime})$, if $c^{\prime}(e)=R$, then $e\in P_{1}$. Otherwise, $e\in P_{2}$.
Observe that by the choice of $c^{\prime}$, neither $G[P_{1}]$ nor $G[P_{2}]$ contain an induced copy of $\textrm{KG}(3r,r)$. Indeed, if $G[P_{i}]$ contains an induced copy of $\textrm{KG}(3r,r)$, then there exists $S\subset[n]$ of size $3r$ such that ${S\choose r}\subset P_{i}$. But this implies that $c^{\prime}$ contains a monochromatic copy of $K^{r}_{3r}$ on vertex set $S$.
@@ 4.2 Proof of Theorem 14
We now define $c:E(G)\to\{R,B\}$ as follows. For any edge $e\in G[P_{R}]$ or $e\in G[P_{B}]$, set $c(e)=R$. Otherwise, set $c(e)=B$. Since the blue subgraph is a bipartite graph, it is triangle-free. Hence $c$ is a red/blue edge coloring of $G$ containing no red induced $\textrm{KG}(3r,r)$ and no blue triangle. ∎
@@ 5 Concluding remark
Appendix A Appendix: Explicit colorings.
A.1 A good coloring of $\textrm{KG}(8,2)$
Here we provide a coloring of $\textrm{KG}(8,2)$ with no monochromatic $K_{3}$. This shows $R^{\textrm{KG}}_{2}(3,3)\geq 9$.
@@ A.1 A good coloring of $\textrm{KG}(8,2)$
Edges from $A$ to $C$:
((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,5),(2,3)),((1,6),(2,3)),((1,7),(2,3)),((1,8),(2,3))((1,2),(3,4)),((1,2),(3,5)),((1,2),(3,6)),((1,2),(3,7)),((1,2),(3,8)),((1,3),(2,4)),\\ ((1,3),(2,5)),((1,3),(2,6)),((1,3),(2,7)),((1,3),(2,8)),((1,4),(2,3)),((1,5),(2,3)),\\ ((1,6),(2,3)),((1,7),(2,3)),((1,8),(2,3))
Edges from $B$ to $C$:
((1,4),(5,6)),((1,4),(6,8)),((1,5),(6,7)),((1,5),(6,8)),((1,5),(7,8)),((1,6),(4,5)),((1,6),(5,7)),((1,6),(5,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(4,6)),((1,8),(4,7)),((1,8),(6,7)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),((2,6),(4,5)),((2,6),(5,7)),((2,6),(5,8)),((2,7),(4,5)),((2,7),(4,8)),((2,7),(5,8)),((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),((3,8),(6,7))((1,4),(5,6)),((1,4),(6,8)),((1,5),(6,7)),((1,5),(6,8)),((1,5),(7,8)),((1,6),(4,5)),\\ ((1,6),(5,7)),((1,6),(5,8)),((1,7),(4,8)),((1,7),(5,8)),((1,8),(4,6)),((1,8),(4,7)),\\ ((1,8),(6,7)),((2,4),(5,6)),((2,4),(6,8)),((2,5),(6,7)),((2,5),(6,8)),((2,5),(7,8)),\\ ((2,6),(4,5)),((2,6),(5,7)),((2,6),(5,8)),((2,7),(4,5)),((2,7),(4,8)),((2,7),(5,8)),\\ ((2,8),(4,5)),((2,8),(4,6)),((2,8),(4,7)),((3,4),(5,6)),((3,4),(6,8)),((3,5),(6,7)),\\ ((3,5),(6,8)),((3,5),(7,8)),((3,6),(4,5)),((3,6),(5,7)),((3,6),(5,8)),((3,7),(4,8)),\\ ((3,7),(5,8)),((3,8),(4,6)),((3,8),(4,7)),((3,8),(6,7))
Edges within $B$:
((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(7,8)),((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),((4,8),(6,7)),((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))((4,5),(6,7)),((4,5),(6,8)),((4,5),(7,8)),((4,6),(5,7)),((4,6),(5,8)),((4,6),(7,8)),\\ ((4,7),(5,6)),((4,7),(5,8)),((4,7),(6,8)),((4,8),(5,6)),((4,8),(5,7)),((4,8),(6,7)),\\ ((5,6),(7,8)),((5,7),(6,8)),((5,8),(6,7))
Edges within $C$:
((1,4),(2,6)),((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,5),(2,6)),((1,5),(2,7)),((1,5),(3,6)),((1,5),(3,7)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),((1,7),(2,5)),((1,7),(2,8)),((1,7),(3,5)),((1,7),(3,8)),((1,8),(2,4)),((1,8),(2,7)),((1,8),(3,4)),((1,8),(3,7)),((2,4),(3,6)),((2,4),(3,8)),((2,5),(3,6)),((2,5),(3,7)),((2,6),(3,4)),((2,6),(3,5)),((2,7),(3,5)),((2,7),(3,8)),((2,8),(3,4)),((2,8),(3,7))((1,4),(2,6)),((1,4),(2,8)),((1,4),(3,6)),((1,4),(3,8)),((1,5),(2,6)),((1,5),(2,7)),\\ ((1,5),(3,6)),((1,5),(3,7)),((1,6),(2,4)),((1,6),(2,5)),((1,6),(3,4)),((1,6),(3,5)),\\ ((1,7),(2,5)),((1,7),(2,8)),((1,7),(3,5)),((1,7),(3,8)),((1,8),(2,4)),((1,8),(2,7)),\\ ((1,8),(3,4)),((1,8),(3,7)),((2,4),(3,6)),((2,4),(3,8)),((2,5),(3,6)),((2,5),(3,7)),\\ ((2,6),(3,4)),((2,6),(3,5)),((2,7),(3,5)),((2,7),(3,8)),((2,8),(3,4)),((2,8),(3,7))